OpenStudy (anonymous):

Log Help: 5^(4x-3)=20 Can anyone help I am not sure how to do this step by step

jimthompson5910 (jim_thompson5910):

In order to isolate the exponent to solve for the variable, you need to use logs (which undo exponents) \[\large 5^{4x-3}=20\] \[\large \log\left(5^{4x-3}\right)=\log(20)\] \[\large (4x-3)\log\left(5\right)=\log(20)\] \[\large 4x*\log\left(5\right)-3\log\left(5\right)=\log(20)\] I'll let you finish up

OpenStudy (anonymous):

Wow thanks.

jimthompson5910 (jim_thompson5910):

np